OpenStudy (anonymous):

Maddy works for the department of soical sercuirty on an annual salary of $52 460. If she recives 17 1/2% holiday loading on the four weeks holiday pay period, calcuate her total holiday pay for the 4 weeks. i need working out please but simply with the answer

OpenStudy (anonymous):

please someone help me

OpenStudy (anonymous):

somebody please

OpenStudy (anonymous):

52,460 is annual salary for 52 weeks. So, weekly pay = 52,460/52. Now, she gets 17.5% extra for the the holiday period of 4 weeks. Multiply weekly pay by 4 and 17.5/100. Then, add this number and the four week pay to get the total number.

OpenStudy (anonymous):

i need to equation i still dont get it

OpenStudy (anonymous):

GT is trying to encourage you to do a bit of the work yourself, so you get an understanding. Work out the weekly wage Maddy earns by dividing the annual salary by 52 (being 52 weeks in a year). Work out what 17.5% of this weekly wage is, and add it to her weekly wage over weeks to figure out what she's earned. It's explained quite well by GT.

OpenStudy (anonymous):

it is 1008.846154 but i still dont get what 17.5% is divided by sorry but yeah

OpenStudy (anonymous):

NORMAL weekly pay = 52,460/52. FOR 4 WEEKS=4*(52,460/52) NOW SHE GETS EXTRA PAY DURING HOLIDAYS thats all so (17.5/100){that is 17.5%} (17.5/100)*4*NORMAL WEEKLY PAY THIS AMNT PLUS NORMAL PAY GIVES THE FINAL INCREASED PAY SHE GETS

OpenStudy (anonymous):

17.5% is 0.175 as a decimal. Simply divide it by 100. A simple way to add an additional x percentage to something is multiply it by 1+x, so in this case 1+0.175 = 1.175. If her normal pay is (52460/52) per week then for the four weeks she receives holiday loading her pay will be: (52460/52)*1.175*4 = $4,741.58.